OpenStudy (ray1998):

Please help! Will give fan and medal. Thank you in advance! :) Find the 16th term of the sequence. a(17) = -2, d = -19 The 16th term of the sequence is what?

OpenStudy (nubeer):

a(17) = a1 + (17-1) * -19 find a1 first

OpenStudy (nubeer):

i hope you are familiar with this formula

OpenStudy (ray1998):

Sorry! Got sidetracked for a minute.

OpenStudy (ray1998):

Is it a1 - a2, or something like that?

OpenStudy (nubeer):

xn = a + d(n-1) this is the general formula for finding any arithmetic sequence

OpenStudy (nubeer):

now plug in the values already according to your question, which i did in the first comment

OpenStudy (ray1998):

is it -19?

OpenStudy (nubeer):

are you asking me for an answer? because honetly i have solved it yet

OpenStudy (ray1998):

No, sorry, I got confused lol. After simplifying though, I have a = a(1) - 304/17

OpenStudy (nubeer):

:/ duh

OpenStudy (nubeer):

lets begin from start ok?

OpenStudy (ray1998):

What?

OpenStudy (ray1998):

Okay.

OpenStudy (nubeer):

General Formula a(n) =a(1) + (n-1)*d got this much?

OpenStudy (ray1998):

Yes

OpenStudy (nubeer):

okay.. in question you are provided with some information for example. a(17)=-2 and d=-19 right?

OpenStudy (ray1998):

Right

OpenStudy (nubeer):

ok plug this information into the general equation.. can u do that?

OpenStudy (nubeer):

so u can find a(1) as its value is missing, and thats one of the key element in finding any arithmetic sequence.

OpenStudy (ray1998):

Okay. I may need some help with that :/ this stuff is really confusing for me.

OpenStudy (nubeer):

okay.. just plug in the data which is given in question okay.. do it like this a(17) = a(1) + (17-1)*d you are told the value of a(17) and d .. right? plug them -2 = a(1) + 16*(-19) got this much?

OpenStudy (ray1998):

Yes

OpenStudy (nubeer):

okay then solve it -2 = a(1) + (-304) solve and find a(1) for me please

OpenStudy (ray1998):

a(1) = 302?

OpenStudy (nubeer):

yes..

OpenStudy (nubeer):

now you have to find a(16) right? so solve it. a(16) = a(1) + (16-1) * d you have a(1) and you have value of d too.. can you do it ?

OpenStudy (ray1998):

Right. Okay.. a(1) = 4 + 16a?

OpenStudy (nubeer):

duh what? how did u got this? a(16) = a(1) + (16-1) * d i wrote this :3

OpenStudy (nubeer):

a(16) = a(1) + (16-1) * d just solve this part.. you just found a(1) above right? just put it in this eauqtion

OpenStudy (ray1998):

I forgot to plug in 302 for a(1), sorry. a = 149/8

OpenStudy (nubeer):

:/ a(16) = 302 + (15*(-19)) a(16) =302 + (-285) a(16) = 17 which part was hard for u :/

OpenStudy (ray1998):

All of it I guess. Math is not my strongest subject, in any sort.

OpenStudy (ray1998):

I apologize for your trouble and time! But thank you, I do appreciate it.

OpenStudy (nubeer):

hmm well its fine.. just look up where u went wrong.. with what i have told u above the method.. with this u can solve any question

OpenStudy (nubeer):

lol i barely answer questions now.. if i didn't wanted to help you i wouldn't had.. so dont apologize .. i wanted to help you so i did :) no worries

OpenStudy (ray1998):

Alrighty :) well thank you very much!

OpenStudy (nubeer):

no problem at all :) bother me anytime u like